OpenStudy (anonymous):

Find parametric equations for the tangent line at the point (cos(3pi/6),sin(3pi/6), (3pi/6)) on the curve x=cost, y= sint, z=t x(t)=? y(t)=? z(t)=?

OpenStudy (amistre64):

just derive each component on its own metir

OpenStudy (amistre64):

merit even

OpenStudy (amistre64):

the derivative of a vector defined function is the tangent to the curve ... like normal

OpenStudy (amistre64):

its when you derive the components of the surface equations that you get a normal

OpenStudy (amistre64):

x=cost; x' = -sin(t) y= sint; y' = cos(t) z = t ; z' = 1

OpenStudy (amistre64):

now just fill in your given point and youve got the slope of the tangent at least

OpenStudy (amistre64):

i take it the given t is 3pi/6 ....

OpenStudy (amistre64):

establish the values of the given point to create the parametric line equations

OpenStudy (anonymous):

so x=-sin(3pi/6) y= cos(3pi/6) z=1

OpenStudy (amistre64):

that would be the vector tangent at that point yes

OpenStudy (amistre64):

to be lazy about it i would express it like this: x =cos(3pi/6) -sin(3pi/6)t y = sin(3pi/6) +cos(3pi/6)t z = 3pi/6 + t

OpenStudy (amistre64):

3pi/6 ... really? whats wrong with pi/2 lol; 90 degress

OpenStudy (amistre64):

x = t y = 1 z = pi/2 + t maybe if i simplified it right

OpenStudy (amistre64):

you agree?

OpenStudy (anonymous):

agree